Find the particular integral of the equation : (D^2-D')z=e^x+y
1
Expert's answer
2020-11-10T13:04:39-0500

Given

"F(D,D')=(D^2-D')z=e^{x+y}"

Note that "F(1,1)=0" ,thus Particular solution is given by


"P.I=\frac{1}{(D^2-D')}e^{x+y}\cdot 1\\
\implies P.I=e^{x+y}\frac{1}{(D+1)^2-(D'+1)}1=e^{x+y}\frac{1}{(D^2+2D-D')}1\\
\implies P.I=e^{x+y}\bigg[\frac{1}{-D'}\bigg(1-\frac{D^2+2D}{D'}\bigg)^{-1}\cdot 1\bigg]\\=-e^{x+y}\frac{1}{D'}(1)=-e^{x+y}y"

Therefore, "P.I=-ye^{x+y}"